A cylindrical container with a diameter of 6 inches and a height of 1.5 feet, what is the height in inches of the container? What is the volume in cubic inches of the container and if the container was filled with 250 cubic inches of sand, what would be the approximate height in inches?
Answer by mananth(16949)

A cylindrical container with a diameter of 6 inches and a height of 1.5 feet, what is the height in inches of the container? What is the volume in cubic inches of the container and if the container was filled with 250 cubic inches of sand, what would be the approximate height in inches?
...
diameter = 6 inches: radius = 3 inches
height = 1.5 feet : =18 inches.
..
Volume V = pi*r^2*h
V= pi*3^2*18
V= 508.9 in^3
...
250=pi*3^2*h
h=250/pi^r^2
h=250/pi*3^2
h= 8.8 inches
..
m.ananth@hotmail.ca